OpenStudy (anonymous):

To travel 161 miles, it takes Sue, riding a moped, 2 hours less time than it takes Doreen to travel 63 miles riding a bicycle. Sue travels 16 miles per hour faster than Doreen. Find the times and rates of both girls.

OpenStudy (anonymous):

solve the system: 161/(v+16)=t-2 63/v=t Doreen time: 9 hours rate: 7miles/hour Sue time: 7 hours rate: 23 miles/hour

OpenStudy (anonymous):

I had it worked out where I have x+16 and x for speeds, and y and y-2 for times. So I go x+16=63/y, and x=161/y-2.... I'm loading up on tutors this week as I clearly need them... thanks a ton